12x+2y=-4
-5x+y=20

Answer by TimothyLamb(4379)

---
12x+2y=-4
-5x+y=20
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = -2
y = 10
---
system matrix
12 2
-5 1
---
inverse of system matrix
0.045454545 -0.090909091
0.22727273 0.54545455
---
determinant of system matrix = 22
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php